OpenStudy (anonymous):

Completing the Square of x2-10x+22=-2

OpenStudy (anonymous):

start with \[x^2-10x=-24\] then take halve of -10 and square it to get \[(x-5)^2=-24+5^2=-24+25=1\]

OpenStudy (anonymous):

once you have \[(x-5)^2=1\] you know \[x-5=1\text { or } x-5=-1\] so \[x=6\text{ or } x=4\]

OpenStudy (anonymous):

x^2-10 x+22 = -2 Subtract 22 from both sides: x^2-10 x = -24 Add 25 to both sides: x^2-10 x+25 = 1 Factor the left hand side: (x-5)^2 = 1 Take the square root of both sides: sqrt(x-5) = 1 Eliminate the absolute value: x-5 = -1 or x-5 = 1 Add 5 to both sides: x = 4 or x-5 = 1 Add 5 to both sides: x = 4 or x = 6

OpenStudy (anonymous):

thank you (: